A child is to be given acetaminophen 120 mg PO q4h. The bottle reads: 160 mg/5 mL. How many mL should the child be given?
Answer #1

Drug prescribed: 120mg

Available: 160mg/5ml

So,

160/5=32

That is

1ml=32mg

0.5ml=16mg

0.25ml=8mg

2ml=64mg

3ml=96mg

3.5ml=112mg.

3.75ml=112+8mg=120mg

So answer= 3.75ml